Ralph, a 40-year-old nurse who earns $80,000 a year, saves 14% of his annual gross income. Assume that Ralph wants to maintain his exact pre-retirement lifestyle. Calculate Ralph’s wage replacement ratio using the top-down approach (round to the nearest %) and using pre-tax dollars.

PLEASE SHOW ALL WORK, IF IN EXCEL PLEASE SHOW ALL EQUATIONS THANKS!
Answer #1

The calculation of wage replacement ratio is shown below:

1 Earnings 2 Less: Savings 80000 =B1*0.14 3 Less: Medicare and social security tax =B1*0.0765 =B1-B2-B3 4 5 Wage replacement

The resultant figures are shown below:

А. 80,000.00 11,200 1 Earnings 2 Less: Savings Less: Medicare and 3 social security tax 4 Wage replacement 5 ratio 6120 62680